# Security modelling for cyber-physical systems: A systematic literature review

### v1

@@ Abstract
Today, cyber-physical systems (CPS) find themselves at the intersection of digital technology and engineering domains, rendering them high-value targets of sophisticated and well-funded cybersecurity threat actors. Prominent cybersecurity attacks on CPS have brought attention to the vulnerability of these systems, and the soft underbelly of critical infrastructure reliant on CPS. Security modelling for CPS is an important mechanism to systematically identify and assess vulnerabilities, threats, and risks throughout system lifecycles, and to ultimately ensure system resilience, safety, and reliability.
This literature review delves into state-of-the-art research in CPS security modelling, encompassing both threat and attack modelling. While these terms are sometimes used interchangeably, they are different concepts. This article elaborates on the differences between threat and attack modelling, examining their implications for CPS security. A systematic search across leading scientific databases yielded 428 articles, from which 15 were selected and categorised into three clusters: those focused on threat modelling methods, attack modelling methods, and literature reviews. Specifically, we sought to examine what security modeling methods exist today, and how they address real-world cybersecurity threats and CPS-specific attacker capabilities throughout the lifecycle of CPS, which typically span longer durations compared to traditional IT systems.
This article also highlights several limitations in existing research, wherein security models adopt simplistic approaches that do not adequately consider the dynamic, multi-layer, multi-path, and multi-agent characteristics of real-world cyber-physical attacks. Against this backdrop, we propose a unified security modelling framework that aligns threat modelling, attack modelling, and security monitoring across the lifecycle of CPS. Last but not least, this article proposes future research directions in this evolving domain.
@@ 1 Introduction
Cyber-physical systems (CPS) are complex systems that integrate physical processes with computational algorithms and network connectivity (Griffor et al. (2017)). These systems are designed to monitor, control, and optimize physical processes in real time, and are typically found in industrial automation, smart infrastructure, healthcare, and transportation domains. Today, CPS find themselves at the intersection of digital technology and engineering, and consequently become a high-value target of cybersecurity threat actors. Prominent cybersecurity attacks on Cyber-Physical Systems (CPS) (Kumar et al. (2022)) have brought attention to the vulnerability of these systems, and the soft underbelly of critical infrastructure reliant on CPS. These vulnerabilities could potentially result in significant consequences, affecting lives, economies, and national security.
@@
Although CPS cybersecurity is often discussed in the context of critical infrastructure, CPS are increasingly interconnected in smart city infrastructure and the cloud. This presents cybersecurity risks not only from physical and cyber components, but also from less conventional but critical attack paths such as electric chargers. For example, Köhler et al. (2023) showed that the Combined Charging System, one of the most widely used DC rapid charging technologies for electric vehicles (EVs), could be exploited wirelessly to interrupt charging sessions for individual vehicles or entire fleets at the same time.
Today, CPS owners and operators face a daunting challenge in keeping pace with the increasing volatility, uncertainty, complexity, and ambiguity (VUCA) of cybersecurity threats. In IT systems, the consequences or impact of IT security incidents are more predictable and can usually be identified and mitigated before actual incidents occur. For example, companies can prepare holding statements to media, and purchase cybersecurity insurance to safeguard against reputational and financial loss before incidents occur. In contrast, CPS cybersecurity attacks are dynamic, less predictable, consequences may change quickly depending on the operating environment, and incident triage and investigations can be especially challenging because of the intricate and interconnected design of CPS.
@@
Existing cybersecurity frameworks like the NIST Cybersecurity Framework (National Institute of Standards and Technology (2023)) which could help organisations manage and improve their cybersecurity risk management, and international standards like IEC 62443-4-1:2018 (International Electrotechnical Commission ) which specifies process requirements for the secure development of products used in industrial automation and control systems, offer valuable guidance for securing CPS. However, the dynamic nature and design constraints of CPS require a tailored approach, combining continual assessment and adaptation of physical measures, cybersecurity practices, as well as risk mitigation controls.
Two key elements of this approach are threat and attack modelling. Although both these terms are often used interchangeably, they are not the same. Threat modelling aims to anticipate cybersecurity threats—usually in the early stages of a system lifecycle, so that relevant mitigation controls can be incorporated into the system design (Uzunov and Fernández (2014); Dhillon (2011); Xiong and Lagerström (2019)). In contrast, attack modelling delves into specific attacker tactics, techniques, and procedures—which may not be limited to cyber but include physical actions as well—to develop targeted defences and mitigation strategies when the system is in operation (Al-Mohannadi et al. (2016); Saini et al. (2008); Schneier (1999); Zenitani (2023)). On this note, threat and attack modelling need to be tailored for CPS and can be an important tool for critical infrastructure owners and operators when navigating cybersecurity threats faced in CPS (Yang and Zhang (2023); Paudel et al. (2017)).
@@
Contributions. In this review, we identified a number of state-of-the-art threat modelling and attack modelling studies relevant to cyber-physical systems (CPS), with the following findings:
@@
  • Threat modelling is commonly conducted in early stages of system development. However, this means that evolving attacker tactics, techniques and procedures (TTPs) in later stages of system lifecycles will render threat models outdated and irrelevant.
@@
  • Security models in the literature generally focus on IT systems, and these models are challenging to use when modelling CPS cybersecurity threats and attacks. This is a pertinent issue for practitioners given the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS.
@@
  • Most papers do not differentiate between cybersecurity breaches in IT systems versus those in CPS. Unlike IT systems, cybersecurity incidents in CPS can result in complex failure modes, as well as consequences in both cyber and physical domains. Adopting a consequence-driven and cyber-informed approach to CPS security is vital towards ensuring that cyber and physical attacks, effects and consequences are considered in security modelling.
@@
  • There is ambiguity in the literature regarding the definitions and relationship between threat modelling and attack modelling. Correspondingly, a unified security modelling framework that integrates threat modelling, attack modelling, and security monitoring to enhance the cyber resilience of CPS is proposed in this article.
@@ 2.1 Cyber-Physical Systems
Despite their longer system lifecycles, CPS are not upgraded or patched frequently as system or software changes may affect their reliability or even safety (Kavallieratos et al. (2020); Kriaa et al. (2015); Sabaliauskaite and Mathur (2014); Suo et al. (2018)). This results in outdated software, exposing CPS to an ever-increasing number of cybersecurity vulnerabilities over time. At the same time, some CPS employ “security-by-obscurity”, using proprietary software and technology that may contain undisclosed security weaknesses that attackers could exploit. Moreover, the lack of “security-by-design” in CPS lowers the resilience to cybersecurity attacks even further, as CPS may lack security logs and fail to leverage log data for proactive cybersecurity monitoring and defence. Finally, there is a misconception that merely physically isolating Cyber-Physical Systems (CPS) from external networks, known as an “air gap” network security measure, is adequate to shield CPS from all cybersecurity threats. However, this notion is flawed, particularly in scenarios where multi-agent, multi-path CPS attacks could employ social engineering techniques, physical access, insider threats, and exploit supply-chain vulnerabilities, enabling attackers to breach the “air gap” and compromise CPS, even when they are physically segregated from external networks.
@@ 2.2 Adversarial model
MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) (Strom et al. (2018)) serves as a widely adopted knowledge base for characterising the actions and behaviours of cyber adversaries. Specifically tailored for Industrial Control Systems (ICS)—a category of Cyber-Physical Systems (CPS)—MITRE ATT&CK provides a matrix known as ”ATT&CK for ICS” (Alexander et al. (2020)). This matrix focuses on the unique Tactics, Techniques, and Procedures (TTPs) employed by adversaries targeting industrial control environments. These TTPs play a crucial role in informing and guiding cybersecurity defence for CPS across the various stages of a cyber-attack lifecycle, as outlined in the Lockheed Martin cyber kill chain model (Hutchins et al. (2011)).
An analysis of past cyber attacks on CPS reveals a taxonomy (Zhu et al. (2011)) highlighting differences from those observed in IT systems, necessitating a tailored security modelling approach for CPS. CPS attackers employ dynamic, multi-agent, and multi-path attack TTPs to defeat physical access controls and “airgap” measures (isolating CPS from external networks). Additionally, adversaries evolve their TTPs over time, such as using “low and slow” TTPs to evade detection and prolong access to compromised CPS.
@@ 2.3 CPS cyber-intrusion example
We present an example of a 6-month long CPS intrusion by a sophisticated attacker group comprising two adversaries A and B, adapted from one of the author’s professional experience. In this example, Adversary A compromises the organisation’s IT network to gain remote access to a CPS engineer’s computer and employs social engineering techniques to deliver custom malware to an isolated (“air gapped”) CPS network. Adversary B’s focus on the other hand is to develop custom malware to firstly profile the CPS, then to weaponise custom malware to manipulate connected Remote Terminal Units (RTUs) causing physical damage e.g. via actuators or valves. The diamond model of intrusion analysis (Caltagirone et al. (2013)), as well as a combination of the diamond model and the cyber kill-chain (Ertaul and Mousa (2018)) are used to describe both the flow (Figure 1) and timeline (Table 1) of this CPS cyber-attack.
Figure 1: CPS cyber intrusion flow
Table 1: CPS cyber intrusion timeline
Phase
January
February
March
April
May
June
Reconnaissance
Scan for email servers, find account credentials
Find users or systems that have access to CPS
Weaponisation
Customise malware for CPS information gathering
Customise malware for RTU reset
Delivery
Deploy malware to CPS through engineer’s USB thumb-drive
Deploy malware to CPS through engineer’s USB thumb-drive
Exploitation
Access VPN
Compromise CPS engineer’s computer
Suppress alarm reporting to main server
Scheduled task manipulates connected RTUs
Installation
Install Cobalt Strike beacon agent
Install Cobalt Strike beacon agent
Malware runs and collects network information
Malware runs on CPS and hides process, awaits scheduled task
C2
Establish C2 callback from IT network
Establish C2 callback
Maintain access
CPS information saved in USB thumb-drive
Maintain access
Maintain access
Actions on Objectives
Services disrupted
@@ 2.4 Security modelling framework for CPS
Figure 2: Security modelling framework for CPS
@@ 3 Related Work
There are few existing systematic literature reviews (SLR) on threat modelling or attack modelling. Humayed et al. (2017)’s literature survey covered the security and privacy of cyber-physical systems, with a special focus on ICS, smart grids, medical devices, and smart cars. They presented a taxonomy of CPS threats, vulnerabilities, known attacks and existing controls, and importantly, captured how an attack of the physical domain of a CPS can result in unexpected consequences in the cyber domain and vice versa.
The SLR by Xiong and Lagerström (2019) was another early review on cybersecurity threat modelling, where one CPS threat modeling method was cited: Burmester et al. (2012)’s framework. Based on the Byzantine paradigm, it enables the modeling of CPS security and facilitates formal analysis and security proof through cryptographic methodologies. Tatam et al. (2021) reviewed threat modelling approaches for APT-style attacks, but the review did not focus on CPS security. The SLR by Khalil et al. (2023) on the other hand, focused on threat modelling of industrial control systems (ICS). The authors described the various ICS threat models adopted in literature, and noted it was timely to consider a framework that covers all aspects of the various threat modelling methods.
Earlier, the evaluation of threat modelling methods by Shevchenko et al. (2018) made a similar conclusion that there is no one method that can cover the full spectrum of CPS threats, and a framework that employed a combination of methods and techniques should be used. Supporting this argument, the hybrid threat model for smart systems proposed by Valenza et al. (2023), and the work by Li et al. (2021) on security-usability threat modelling for industrial control systems underscored the need to consider not only cyber but also human factors in ICS threat modelling due to the system-of-systems (SoS) nature of CPS.
These CPS threat modelling methods and techniques were well-covered in the literature, with case studies in specific industries and domains, e.g. the study of smart grid CPS cybersecurity by Nafees et al. (2023), the study of cyber-physical energy systems security (CPES) by Zografopoulos et al. (2021), the application of the STRIDE threat model in modern vehicle vulnerability assessments by Abuabed et al. (2023), the work on CPS threat modelling for power transformers by Ahn et al. (2021), the work on medical CPS threat modelling by Almohri et al. (2017), the work by Al Asif and Khondoker (2020) on telesurgery system threat modelling, the work by Jbair et al. (2022) on threat modelling for industrial cyber-physical systems (ICPS) in the manufacturing industry, and the work on threat modelling in smart firefighting systems by Zahid et al. (2023).
@@ 4 Methodology
We followed the guidelines for systematic literature reviews proposed by Kitchenham and Charters (2007). The process of our literature review is shown in Figure 3.
Figure 3: Literature review process
@@ 4.1 Research questions
  • RQ1: What cyber-physical system (CPS) threat or attack models have been adopted in existing literature?
@@
  • RQ1.1: How do the proposed threat and attack models align with cybersecurity in various phases of the cyber-physical system lifecycle?
@@
  • RQ2: What research gaps have been identified in the methods used for modelling threats and attacks in cyber-physical systems (CPS)?
@@ 4.2 Search protocol
This sub-section aims to describe the search protocol that we used in this review.
We used US English for all search terms. For example, we used “modeling” instead of “modelling”, in UK English. We validated that Scopus did not distinguish between the spellings and returned the same search results (Figure 4).
Figure 4: Comparison of search results between US and UK English keywords
We selected IEEE Computer Society Digital Library, Science Direct, ACM Digital Library and Scopus to perform our search for articles. These scientific databases were chosen as they are recognised as reputable and accurate repositories for academic literature, and they have a good collection of results relating to cybersecurity articles and publications.
For the search keywords, we were primarily interested to search: cybersecurity modelling, security modelling, threat modelling and attack modelling, in line with the subject of our literature review. As such, we specified our first search keywords as: “security model”, “security modeling”, “cybersecurity model”, “cybersecurity modeling”, “threat model”, “threat modeling”, “attack model”, and “attack modeling”. We included both words, “model” and “modeling” to increase the coverage of search results.
As we wanted to narrow our search results to articles on security, threat, or attack modelling, we limited the initial search to article titles. There was a need to make minor changes to the search terms, depending on the Boolean vocabulary used on the database.
  • On IEEE Computer Society Digital Library, we selected Advanced Search, selected the “in Document Title” option, and entered the search keywords: (model modeling) AND (cybersecurity security threat attack).
@@
  • On ACM Digital Library, we selected “Edit Search”, selected “Title” under “Search Within”, then entered: (“security model” “security modeling” “cybersecurity model” “cybersecurity modeling” “threat model” “threat modeling” “attack model” “attack modeling”).
  • On Scopus, we selected “Article title” under the “Search within” field, then entered “security model” OR “security modeling” OR “threat model” OR “threat modeling” OR “attack model” OR “attack modeling” OR “cybersecurity model” OR “cybersecurity modeling” in the “Search documents” field.
We then refined the search results further by specifying additional keywords: (“cyber physical” CPS SCADA ICS APT), to search within article titles, abstracts and contents. We included SCADA (Supervisory Control and Data Acquisition), ICS (Industrial Control System) and APT (Advanced Persistent Threat) in the search, so that search results would be relevant to the adversarial tactics, techniques and procedures associated with CPS cyber attacks.
  • On IEEE Computer Society Digital Library, we used the Advanced Search menu with the previous keywords, selected the “AND” and “in Anything” options, and entered the keywords: (“cyber physical” CPS SCADA ICS APT).
  • On ScienceDirect, we selected Advanced Search and entered (“cyber physical” OR CPS OR SCADA OR ICS OR APT) in the “Find articles with these terms” field.
  • On ACM Digital Library, we selected “Edit Search”, added a new search field, selected “Anywhere” under “Search Within”, then entered: (“cyber physical” OR CPS OR SCADA OR ICS OR APT).
  • On Scopus, we selected “Add search field”, selected “Article title, Abstract, Keywords” under the “Search within” field, and entered: “cyber physical” OR CPS OR SCADA OR ICS OR APT.
In the final phase of the search protocol, we removed repeated articles from the combined search results, based on similar article titles. We also removed articles that were not related to the research questions, after reviewing article titles and abstracts.

### v2

@@ Abstract
Cyber-physical systems (CPS) are at the intersection of digital technology and engineering domains, rendering them high-value targets of sophisticated and well-funded cybersecurity threat actors. Prominent cybersecurity attacks on CPS have brought attention to the vulnerability of these systems, and the inherent weaknesses of critical infrastructure reliant on CPS. Security modelling for CPS is an important mechanism to systematically identify and assess vulnerabilities, threats, and risks throughout system lifecycles, and to ultimately ensure system resilience, safety, and reliability. This survey delves into state-of-the-art research in CPS security modelling, encompassing both threat and attack modelling. While these terms are sometimes used interchangeably, they are different concepts. This article elaborates on the differences between threat and attack modelling, examining their implications for CPS security. A systematic search yielded 428 articles, from which 15 were selected and categorised into three clusters: those focused on threat modelling methods, attack modelling methods, and literature reviews. Specifically, we sought to examine what security modelling methods exist today, and how they address real-world cybersecurity threats and CPS-specific attacker capabilities throughout the lifecycle of CPS, which typically span longer durations compared to traditional IT systems. This article also highlights several limitations in existing research, wherein security models adopt simplistic approaches that do not adequately consider the dynamic, multi-layer, multi-path, and multi-agent characteristics of real-world cyber-physical attacks.
@@ 1 Introduction
Cyber-physical systems (CPS) [1] are facing increasingly unconventional cybersecurity threats. Unlike traditional cybersecurity threats, modern cybersecurity threats are volatile, uncertain, complex, and ambiguous (VUCA). These threats are exacerbated by the wide interconnectivity of CPS in smart city infrastructure and the cloud. This interconnectivity presents a myriad of cybersecurity risks from both physical and cyber domains, including unconventional attack paths. In one example of an unconventional attack path, Kohler et al. [2] demonstrated that the Combined Charging System, a widely used DC rapid charging technology for electric vehicles (EVs), could be exploited wirelessly to interrupt charging sessions for individual vehicles or entire fleets at the same time.
The severity of CPS cyber-attacks stem from the roles that CPS perform in monitoring, controlling, and optimising critical operations that serve cities and communities. The critical role of CPS further underscores the need to ensure they are resilient to cyber attacks. However, consequences of CPS cybersecurity attacks can be less predictable and far more difficult to mitigate if cybersecurity threats were not adequately considered at the point of system design.
CPS owners and operators face a daunting challenge in keeping pace with the modern cybersecurity landscape. They find themselves at the intersection of digital technology and engineering and are often targeted in sophisticated cybersecurity attacks. Prominent cybersecurity attacks [3] on CPS have brought attention not only to the vulnerability of these systems, but also the inherent weaknesses of critical infrastructure reliant on them. These vulnerabilities could potentially result in severe consequences, affecting lives, economies, and national security.
@@
To address these challenges, CPS owners and operators can take guidance from internationally recognised cybersecurity frameworks and standards. Firstly, the National Institute of Standards and Technology (NIST) Cybersecurity Framework [4] provides guidance to identify and manage cybersecurity risks in a systematic way. Secondly, the International Electrotechnical Commission (IEC) 62443-4-1:2018 standard [5] specifies process requirements for the secure development of products used in industrial automation and control systems, offering valuable guidance for securing CPS. These alone are not sufficient to mitigate ever-evolving cybersecurity challenges facing CPS. The dynamic threat landscape requires a tailored approach to CPS cybersecurity, combining continual assessment and adaptation of physical measures, cybersecurity practices, as well as risk mitigation controls.
To establish a clear framework for this paper, we define three key terms: “security modelling”, “threat modelling”,and “attack modelling”. Security modelling refers to the overall process of identifying, assessing, and mitigating risks to protect a system or organisation from a wide range of potential threats, both cyber and physical, throughout its entire lifecycle. Within this broader framework, threat modelling is a proactive approach used primarily in the early stages of a system’s lifecycle to anticipate potential cybersecurity threats and vulnerabilities, enabling the incorporation of relevant mitigation controls into the system design [6] [7] [8]. In contrast, attack modelling is a reactive approach that focuses on analysing specific attacker tactics, techniques, and procedures—including both cyber and physical actions—once the system is in operation, to develop targeted defenses and response strategies [9] [10] [11] [12]. To effectively mitigate risks in CPS, integrating threat and attack modelling into a continuous, iterative security modelling process is essential throughout the system’s lifecycle [13] [14].
In this paper, we design and conduct a systematic literature review (SLR) to study how existing security modelling, threat modelling and attack modelling approaches may be integrated in an iterative framework tailored to CPS in a continuous lifecycle context. This framework is crucial to ensure that CPS security measures are not static but evolve with the changing cybersecurity threat landscape. By bridging this gap, our survey aims to propose a holistic approach to enhance the cyber resilience of CPS.
@@
Findings. In this review, we identified several state-of-the-art security modelling, threat modelling and attack modelling studies relevant to CPS, with the following findings:
@@
  • Threat modelling is usually conducted in early stages of system development, and unlikely to anticipate new attacker tactics, techniques and procedures (TTPs) in later stages of system lifecycles.
@@
  • Security models in the literature that focus on IT systems could be challenging to use when modelling CPS cybersecurity threats and attacks. This is a pertinent issue for practitioners given the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS.
@@
  • There is limited differentiation between cybersecurity breaches in IT systems versus those in CPS. Unlike IT systems, cybersecurity incidents in CPS can result in complex failure modes, as well as consequences in both cyber and physical domains. Adopting a consequence-driven and cyber-informed approach to CPS security is vital towards ensuring that cyber and physical attacks and consequences are considered in security modelling.
@@
  • Unclear definitions and relationships between threat modelling and attack modelling can lead to inconsistent security modelling approaches. A unified security modelling approach that integrates threat modelling, attack modelling, and security monitoring to enhance the cyber resilience of CPS is proposed in this article to address this.
@@ 2.1 Cyber-Physical Systems
Despite their longer system lifecycles, CPS are not upgraded or patched frequently as system or software changes may affect their reliability or even safety [15] [16] [17] [18]. This results in outdated software, exposing CPS to an ever-increasing number of cybersecurity vulnerabilities over time. At the same time, some CPS employ “security-by-obscurity”, using proprietary software and technology that may contain undisclosed security weaknesses that attackers could exploit. Moreover, the lack of “security-by-design” in CPS lowers the resilience to cybersecurity attacks even further, as CPS may lack security logs and fail to leverage log data for proactive cybersecurity monitoring and defence. Finally, there is a misconception that merely physically isolating CPS from external networks, known as an ”air gap” network security measure, is adequate to shield CPS from all cybersecurity threats. However, this notion is flawed, particularly in scenarios where multi-agent, multi-path CPS attacks could employ social engineering techniques, physical access, insider threats, and exploit supply-chain vulnerabilities, enabling attackers to breach the ”air gap” and compromise CPS, even when they are physically segregated from external networks.
@@ 2.2 Adversarial model
MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) [19] serves as a widely adopted knowledge base for characterising the actions and behaviours of cyber adversaries. Specifically tailored for ICS—a category of CPS—MITRE ATT&CK provides a matrix known as ”ATT&CK for ICS” [20]. This matrix focuses on the unique Tactics, Techniques, and Procedures (TTPs) employed by adversaries targeting industrial control environments. These TTPs play a crucial role in informing and guiding cybersecurity defence for CPS across the various stages of a cyber-attack lifecycle, as outlined in the Lockheed Martin cyber kill chain model [21].
An analysis of past cyber attacks on CPS reveals a taxonomy [22] highlighting differences from those observed in IT systems, necessitating a tailored security modelling approach for CPS. CPS attackers employ dynamic, multi-agent, and multi-path attack TTPs to defeat physical access controls and ”airgap” measures (isolating CPS from external networks). Additionally, adversaries evolve their TTPs over time, such as using ”low and slow” TTPs to evade detection and prolong access to compromised CPS.
@@ 2.3 CPS cyber-intrusion example
We present an example of a 6-month long CPS intrusion by a sophisticated attacker group comprising two adversaries A and B, adapted from one of the author’s professional experience. In this example, Adversary A compromises the organisation’s IT network to gain remote access to a CPS engineer’s computer and employs social engineering techniques to deliver custom malware to an isolated (”air gapped”) CPS network. Adversary B’s focus on the other hand is to develop custom malware to firstly profile the CPS, then to weaponise custom malware to manipulate connected Remote Terminal Units (RTUs) causing physical damage e.g. via actuators or valves. The diamond model of intrusion analysis [23] as well as a combination of the diamond model and the cyber kill-chain [24] are used to describe both the flow (Figure 1) and timeline (Figure 2) of this CPS cyber-attack.
Figure 1: CPS cyber intrusion flow representation using the diamond model of intrusion analysis
Figure 2: Timeline of a CPS cyber intrusion illustrating the steps in a cyber kill-chain
@@ 2.4 Security modelling framework for CPS
Figure 3: Security modelling framework for CPS that integrates threat modelling, attack modelling and security monitoring throughout the system lifecycle
@@ 3 Related Work
There are few existing systematic literature reviews on threat modelling or attack modelling. Humayed et al.’s literature survey [25] covered the security and privacy of cyber-physical systems, with a special focus on ICS, smart grids, medical devices, and smart cars. They presented a taxonomy of CPS threats, vulnerabilities, known attacks and existing controls, and importantly, captured how an attack of the physical domain of a CPS can result in unexpected consequences in the cyber domain and vice versa.
Xiong et al.’s SLR [8] was another early review on cybersecurity threat modelling, where one CPS threat modeling method was cited: Burmester et al.’s framework [26]. Based on the Byzantine paradigm, it enables the modeling of CPS security and facilitates formal analysis and security proof through cryptographic methodologies. Tatam et al. [27] reviewed threat modelling approaches for APT-style attacks, but the review did not focus on CPS security. The SLR by Khalid et al. [28] on the other hand, focused on threat modelling of industrial control systems (ICS). The authors described the various ICS threat models adopted in literature, and noted it was timely to consider a framework that covers all aspects of the various threat modelling methods.
Earlier, the evaluation of threat modelling methods by Shevchenko et al. [29] made a similar conclusion that there is no single method that can cover the full spectrum of CPS threats, and a framework that employed a combination of methods and techniques should be used. Supporting this argument, the hybrid threat model for smart systems proposed by Valenza et al. [30], and the work by Li et al. [31] on security-usability threat modelling for industrial control systems underscored the need to consider not only cyber but also human factors in ICS threat modelling due to the system-of-systems (SoS) nature of CPS.
These CPS threat modelling methods and techniques were well-covered in the literature, with case studies in specific industries and domains, e.g. the study of smart grid CPS cybersecurity by Nafees et al. [32], the study of cyber-physical energy systems security (CPES) by Zografopoulos et al. [33], the application of the STRIDE threat model in modern vehicle vulnerability assessments by Abuabed et al. [34], the work on CPS threat modelling for power transformers by Ahn et al. [35], the work on medical CPS threat modelling by Almohri et al. [36], the work by Al Asif et al. [37] on telesurgery system threat modelling, the work by Jbair et al. [38] on threat modelling for industrial cyber-physical systems (ICPS) in the manufacturing industry, and the work on threat modelling in smart firefighting systems by Zahid et al. [39].
@@ 4 Methodology
We followed the guidelines for systematic literature reviews proposed by Kitchenham et al. [40]. The process of our literature review is shown in Figure 4.
Figure 4: Systematic literature review process
@@ 4.1 Research questions
  • RQ1: What CPS threat or attack models have been adopted in existing literature?
@@
  • RQ1.1: How do the proposed threat and attack models align with cybersecurity in various phases of the CPS lifecycle?
@@
  • RQ2: What research gaps have been identified in the methods used for modelling threats and attacks in CPS?
@@ 4.2 Search protocol
This sub-section outlines the search protocol used in this review.
We used both US and UK English variations for all search terms. For example, we included both ”modelling” (UK English) and ”modeling” (US English).
The databases selected for our search were the IEEE Computer Society Digital Library, Science Direct, ACM Digital Library, and Scopus. These databases were chosen because they are recognised as reputable and reliable repositories for academic literature and have an extensive collection of articles and publications related to cybersecurity.
While multidisciplinary approaches that integrate principles from control systems engineering, safety engineering, and human factors are vital for addressing the intricate interplay between cyber and physical components in CPS environments, they fall outside the scope of this paper. A comprehensive review of the full range of strategies required for effective CPS protection would necessitate insights from various specialised domains, each deserving dedicated attention and expertise.
Our primary focus was therefore on search terms related to cybersecurity modelling, security modelling, threat modelling, and attack modelling, in alignment with the subject of our literature review. We defined Criteria A as our initial list of search keywords, which included: ”security model,” ”security modelling,” ”cybersecurity model,” ”cybersecurity modelling,” ”threat model,” ”threat modelling,” ”attack model,” and ”attack modelling.” We incorporated both ”model” and ”modelling” to maximise the coverage of relevant search results.
As we wanted to narrow our search results to articles on security, threat, or attack modelling, we limited the initial search to article titles. Minor adjustments to the search terms were necessary, depending on the Boolean vocabulary supported by each database.
  • At the time of writing, the IEEE Computer Society Digital Library does not support in-line Boolean operators like ”OR” and ”AND” in the search term input. Therefore, we used the Advanced Search menu, selected “in Document Title,” and conducted repeated searches using individual search keywords, such as ”security modelling” and ”attack model”.
@@
  • On ACM Digital Library, we selected “Advanced Search”, selected “Title” under “Search Within”, then entered: ”security model” OR ”security modeling” OR ”security modelling” OR ”cybersecurity model” OR ”cybersecurity modeling” OR ”cybersecurity modelling” OR ”threat model” OR ”threat modeling” OR ”threat modelling” OR ”attack model” OR ”attack modeling” OR ”attack modelling”.
  • On Scopus, we selected “Article title” under the “Search within” field, then entered ”security model” OR ”security modeling” OR ”security modelling” OR ”cybersecurity model” OR ”cybersecurity modeling” OR ”cybersecurity modelling” OR ”threat model” OR ”threat modeling” OR ”threat modelling” OR ”attack model” OR ”attack modeling” OR ”attack modelling” in the “Search documents” field.
We then refined the search results further by specifying additional keywords: (”cyber physical,” CPS, SCADA, ICS, APT), to search within article contents. We included SCADA (Supervisory Control and Data Acquisition), ICS (Industrial Control System), and APT (Advanced Persistent Threat) to ensure that the search results were relevant to the adversarial tactics, techniques, and procedures associated with CPS cyber attacks. This list of search terms constitutes Criteria B.
  • On the IEEE Computer Society Digital Library, we used the Advanced Search menu with the initial keywords, selected the “AND” and “in Document Content” options, and entered the additional keywords. We noted that there might be duplicate search results due to the use of different additional keywords, which needed to be identified and removed.
  • On ScienceDirect, we selected Advanced Search and entered (”cyber physical” OR CPS OR SCADA OR ICS OR APT) in the “Find articles with these terms” field.
  • On ACM Digital Library, we selected “Edit Search”, added a new search field, selected “Full text” under “Search Within”, then entered: ”cyber physical” OR CPS OR SCADA OR ICS OR APT.
  • On Scopus, we selected “Add search field”, selected “Article title, Abstract, Keywords” under the “Search within” field, and entered: ”cyber physical” OR CPS OR SCADA OR ICS OR APT.
In the final phase of the search protocol, we removed repeated articles from the combined search results based on similar article titles. We also excluded articles that were not related to the research questions, after reviewing their titles and abstracts.